Disk VSN: MED-0040 | Data volume at beginning: 0.0 GB
Note: 0.5 second offset between mark5b and dbbc, as long as the delay is stable we may ignore the alarm. Auto corr looks bad. Warren says it might be a hardware problem. No tsys readings?
1830UT Experiment started okay (Lim).
22:30 UT steady delay offset may be OK but auto correlations are no good, dbbc bad. Power off dbbc and reconfigure. (warren)
22:47 UT complete power off / power on cycle of dbbc seems to have done the trick. first good scan 2247 (w)
23:13 UT halt agian for power cycle and reconfigure to fix two bad channels (w)
23:32 UT restart has made it worse again….(w)
00:24 UT trying a scan, clkoff still bad, autocorrelations bad (w)
01:29 UT still can't get dbbc configured, investigating (w)
02:22 UT back on schedule, dbbc powered off for some time, has cooled? recorded data good from here (w)
13:00 UT find autocorrelations are no good at start of shift, dbbc bad. No idea when it went bad in previous shift. Power off dbbc and reconfigure. still bad. Warren said to keep power cycling and reconfigure dbbc. still working on it. (Jay)
14:30 UT autocorrelations Still bad (Jay)
15:20 UT Tried reconfiguring dbbc several times with giving it time to cool down. clkoff values and autocorr spectra is still bad (Jay).
/home/www/auscope/opswiki/data/pages/handover/r4957.txt · Last modified: 2020/07/31 18:37 by Jayender Kumar